OpenStudy (anonymous):

Luis ran the same 400 m track in 297 strides. Find his stride length in inches.

OpenStudy (mathstudent55):

How many cm to an inch? How many cm to a m?

OpenStudy (anonymous):

I'm terrible at measurements.

OpenStudy (mathstudent55):

Here are the conversions: \(100 ~cm = 1 ~m \) \(2.54 ~cm = 1~in. \)

OpenStudy (mathstudent55):

That means that you cna write the conversion factors like this: \( \dfrac{100 ~cm}{1 ~m} = \dfrac{1~m}{100 ~cm} = 1\) \( \dfrac{2.54 ~cm}{1 ~in.} = \dfrac{1~in.}{2.54 ~cm} = 1\)

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (mathstudent55):

Let's say you need to convert from cm to m. How do you know which conversion factor to use between \( \dfrac{100 ~cm}{1 ~m} = \dfrac{1 ~m}{100 ~cm} \) ? It's simple. You need to use the conversion factor that will cancel out the unit you don't want and will leave the unit you want. For example: Convert 500 cm to m. \( 500 ~cm \times \dfrac{1 ~m}{100 ~cm} = \dfrac{500}{100} ~m = 5 ~m\) Notice that by choosing the conversion factor with cm in the denominator, the cm of the given number cancel out and you are left with just m, which is what the question asked for.